OpenStudy (anonymous):

solve the equation. -9d - 24 = -3(d+8) - 6d

OpenStudy (anonymous):

d is 1

OpenStudy (anonymous):

the 24 cancel out and -9 divided by -9 is 1 so D=1

OpenStudy (anonymous):

-9d-24=-3d-24-6d taking all d on one side and others on another -9d+9d=-24+24 that 0=0 this is true for all values of d so d can be anything

OpenStudy (anonymous):